OpenStudy (anonymous):

7th grade help please

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 4 – 2x + 5x when x = 4 A. –24 B. –18 C. 16 D. 28

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thanks i have more please

OpenStudy (anonymous):

@bryson15 No direct answers please. :)

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 3 – 5z + 4(z – 2) when z = 12 A. –14 B. –98 C. –16 D. –96

OpenStudy (anonymous):

number 1 4 ? 2x + 5x when x = 4 4 - 2(4) + 5(4) = 4 - 8 + 20 = 16

OpenStudy (anonymous):

those question marks r negative signs

OpenStudy (anonymous):

thanks one second

OpenStudy (anonymous):

\[\frac{ k-2 }{ 12\div k}\] when k = −4?

OpenStudy (anonymous):

z ? 3 ? 5z + 4(z ? 2), when z = 12, 12 ? 3 - 5(12) + 4(12 - 2), 4 - 60 + 4(10), 4 - 60 + 40, 4 - 20 = -16 The answer is D. -16

OpenStudy (anonymous):

That would be \(\Huge\frac{-4 - 2}{12 / -4}\)

OpenStudy (anonymous):

is the answer 2?

OpenStudy (anonymous):

@bryson15 Please no direct answers.

OpenStudy (anonymous):

z ? 3 ? 5z + 4(z ? 2), when z = 12, 12 ? 3 - 5(12) + 4(12 - 2), 4 - 60 + 4(10), 4 - 60 + 40, 4 - 20 = -16 The answer is D. -16

OpenStudy (anonymous):

ok i need help with wht i just put up there:)

OpenStudy (anonymous):

thanx bye:D